# Online Multi-Task Offloading for Semantic-Aware Edge Computing Systems

### v1

@@ Abstract
Mobile edge computing (MEC) provides low-latency offloading solutions for computationally intensive tasks, effectively improving the computing efficiency and battery life of mobile devices. However, for data-intensive tasks or scenarios with limited uplink bandwidth, network congestion might occur due to massive simultaneous offloading nodes, increasing transmission latency and affecting task performance. In this paper, we propose a semantic-aware multi-modal task offloading framework to address the challenges posed by limited uplink bandwidth. By introducing a semantic extraction factor, we balance the relationship among transmission latency, computation energy consumption, and task performance. To measure the offloading performance of multi-modal tasks, we design a unified and fair quality of experience (QoE) metric that includes execution latency, energy consumption, and task performance. Lastly, we formulate the optimization problem as a Markov decision process (MDP) and exploit the multi-agent proximal policy optimization (MAPPO) reinforcement learning algorithm to jointly optimize the semantic extraction factor, communication resources, and computing resources to maximize overall QoE. Experimental results show that the proposed method achieves a reduction in execution latency and energy consumption of 18.1% and 12.9%, respectively compared with the semantic-unaware approach. Moreover, the proposed approach can be easily extended to models with different user preferences.
@@ I Introduction
Nevertheless, when tasks involve substantial data transfers, such as high-definition (HD) video processing [2], virtual reality (VR) applications [3], or real-time data analytics [4], the increased volume of data transmission imposes higher demands on channel quality to meet latency requirements. In such cases, semantic communication (SemCom) can provide an effective solution. SemCom, different from the traditional bit workflow based on Shannon’s information theory [5], has demonstrated great potential for enhancing the efficiency and reliability of data transmission across various task scenarios, including text [6, 7, 8], image [9, 10, 11], video [12, 13, 14], XR [15, 16, 17], and multi-modal tasks [18, 19, 20]. By transmitting a compact semantic representation of task data, substantial bandwidth can be saved, ensuring high quality of service (QoS).
@@ I-A Related Works
Given these considerations, it is essential to design a semantic-aware MEC system by combining MEC with SemCom [30]. By performing semantic extraction on the data to be offloaded, UEs can filter out data irrelevant to the task, thereby reducing network traffic. In [31], the authors defined a semantic entropy to quantify the semantic information of different tasks. Deep Q-Network (DQN) and many-to-one matching algorithms are used to solve the joint optimization problem of semantic compression, channel allocation, and power distribution in multi-cell multi-task scenarios. A semantic-aware resource allocation method was proposed in [32], introducing a semantic-aware factor to characterize the relationship between semantic compression rate and computational overhead, aiming to balance the degree of semantic compression. The authors of [33] proposed a semantic-aware cloud-edge-end collaborative networking method to utilize distributed computing resources and also proposed a multi-modal task offloading system in [34], where a unified QoE metric was designed.
@@ I-B Motivations and Contributions
In traditional MEC scenarios, computational and communication resources are often jointly optimized to minimize system energy consumption [35] or execution latency [36], or a combination of both [37]. However, these approaches lack consideration for task execution performance and comprehensive QoE. 6G strives to provide stable, reliable, and high-quality personalized services to each user [38]. Preferences often vary among different users and types of services. For instance, in MEC scenarios, users dealing with latency-sensitive tasks may prefer lower latency, those pursuing task performance may favor better task execution accuracy, while users with limited battery power might opt for task offloading strategies that consume less power. Therefore, a good QoE metric should encompass the preferences of different users. Additionally, since each user may need to handle multi-modal tasks, the complexity and evaluation metrics of different tasks vary. For example, bimodal tasks like visual question answering (VQA) typically require more processing time and computational resources compared to text translation, and their evaluation metrics are entirely different.
Semantic-aware MEC systems have been studied in [31, 32, 33, 34]. Nonetheless, In [31, 32, 33], the authors did not address multi-modal tasks or design a unified QoE metric. The authors of [34] did not consider the impact of semantic extraction on task execution latency and performance. To address the aforementioned challenges, we investigate resource allocation schemes in multi-modal task offloading that coordinate computational and communication resources through semantic compression-aware methods. To fairly compare the offloading performance of different task modalities, a unified QoE metric is proposed. Computational and communication resources are jointly optimized to maximize the overall QoE for UEs. The main contributions of this paper are summarized as follows:
  • We propose a novel semantic-aware task offloading system for multi-modal tasks. The SemCom systems for image classification, text classification, and VQA are first constructed to compress source data into compact semantic representations. The semantic extraction factor is used to characterize the degree of data compression, which affects the task’s execution latency, energy consumption, and performance.
@@
  • The joint optimization for semantic extraction factor, computational resources, and communication resources is performed using the MAPPO algorithm, maximizing the overall QoE. Users solve for the optimal channel allocation, power allocation, semantic extraction level, and computation offloading scheme according to their observed task sizes and channel conditions.
  • Comprehensive experiments demonstrate the effectiveness of the proposed approach. Compared to the semantic-unaware method, the proposed algorithm achieves a reduction in execution latency and energy consumption of 18.1% and 12.9%, respectively. Additionally, the proposed approach can meet user preferences for various metrics such as execution latency, energy consumption, and execution performance.
@@ II System Model and Problem Formulation
We propose a novel semantic-aware multi-modal task offloading system, as shown in Fig. 1, where UEs employ an orthogonal frequency-division multiple access (OFDMA) scheme. At the access point (AP), $K$ resource blocks are allocated to serve $N$ users at different sub-channels. The set of UEs and sub-channels are defined as $\mathcal{N}=\{1,2,\ldots,N\}$ and $\mathcal{K}=\{1,2,\ldots,K\}$. The length of the UEn’s task queue is denoted as $Q_{n}$. The task currently being executed by the UEn is represented by $q^{\mathcal{M}}_{n}$, where $\mathcal{M}\in\{\mathcal{T}:text,\mathcal{I}:image,\mathcal{B}:VQA\}$. Hence, each UE is associated with three potential task types: text classification, image classification, and VQA. We assume that tasks are atomic [34], i.e. all tasks are either executed locally or offloaded to the ES. The illustrations of the main notations are summarized in Table I.
@@ II-A Semantic-Aware Model
Traditional MEC offloading methods usually require transmitting all data to the ES. SemCom enables the extraction of task-relevant semantic information, reducing bandwidth consumption while enhancing robustness to channel variations. Based on this, we propose a semantic-aware multi-modal task offloading framework. Fig. 2 illustrates the process of semantic-aware offloading. While semantic extraction indeed reduces the volume of data transmitted, the associated increase in semantic extraction latency and computational burden must also be comprehensively considered. Hence, we define a semantic extraction factor $\mu_{n}\in\left[\mu_{n}^{min},1\right]$. The semantic extraction module determines the level of data compression. When $\mu_{n}=1$, the UE transmits all data instead of semantic information. Transmitting all data can achieve better task performance but imposes a heavy transmission burden. The smaller the $\mu_{n}$, the greater the level of semantic compression, which significantly reduces the transmission load. However, semantic extraction introduces additional computational overhead and may decrease task performance on the ES. The minimum value $\mu_{n}^{min}$ is typically determined by task accuracy requirements and user battery capacity. Therefore, selecting $\mu_{n}$ requires a comprehensive evaluation of the current task computational load, channel environment, user battery capacity, and other relevant information.
Figure 1: Semantic-aware multi-task offloading system model.
@@ II-B Transmission Model
The channel gain from UEn to ES via sub-channel $k$ is denoted by $|g_{k,n}|^{2}=|h_{k,n}|^{2}/d_{n}^{-\alpha}$, where $h_{n}\sim\mathcal{C}\mathcal{N}(0,1)$ is Rayleigh fading coefficient, and $d_{n}$ is the distance from UEn to ES, and $\alpha$ is the path loss exponent, and $\mathcal{C}\mathcal{N}(0,1)$ is the complexed Gaussian distribution with zero mean and variance one. The transfer function can be represented as
$y_{n}=g_{k,n}\sqrt{p_{n}}s_{n}+n_{z},$ (1)
where $p_{n}$ denotes the transmit power from UEn to ES, and $s_{n}$ is the normalized transmission data from UEn to ES, and $n_{z}\sim\mathcal{C}\mathcal{N}(0,\sigma_{z}^{2})$ is zero-mean additive white Gaussian noise (AWGN) with noise power $\sigma_{z}^{2}$. The maximum achievable offloading data rate from UEn to ES can be represented as
$R_{k,n}=B\text{log}_{2}\left(1+\frac{g_{k,n}p_{n}}{\sigma_{z}^{2}}\right),$ (2)
where $B$ is the channel bandwidth. Therefore, the transmission latency can be expressed as
$t_{n}=\frac{\rho_{n}s_{n}\mu_{n}^{p}}{R_{k,n}},$ (3)
where $p>0$ is a constant parameter. $\rho_{n}$ is a binary variable, where $\rho_{n}=1$ represents offloading tasks to ES, and $\rho_{n}=0$ denotes local computation. In comparison to the amount of data required for task offloading, the computation results typically involve a much smaller data size [39, 40], therefore, we only consider the upload rate and latency.
@@
$\mathcal{M}$ The set of task modalities
@@ II-C Computation Model
The computation latency of UEn can be denoted as
$t_{n}^{U}=\begin{cases}\frac{(1-\rho_{n})l_{n}^{U}}{c_{n}},\ \text{if}\ \mu_{n}=1,\\ \frac{(1-\rho_{n})l_{n}^{U}}{c_{n}}+\frac{\rho_{n}l_{n}^{SE}}{\mu_{n}^{k}c_{n}},\ \text{otherwise},\end{cases}$ (4)
where $l_{n}^{U}$ is the Graphic Processing Unit (GPU) consumption of the current task computed by UEn, and $l_{n}^{SE}$ is the basic GPU consumption of semantic extraction. $l_{n}^{U}$ and $l_{n}^{SE}$ are related to the corresponding task $\mathcal{M}$. $k>0$ is a constant parameter. $c_{n}=o_{n}f_{n}$ is the GPU computation capability of UEn. $o_{n}$ represents the floating-point operations (FLOPS) executed per cycle by the GPU of UEn and $f_{n}$ denotes the GPU clock frequency of UEn. When tasks are offloaded to ES, the computation time of tasks on the ES can be expressed as
$t_{n}^{S}=\frac{l_{n}^{S}c_{s}}{\sum_{n\in\mathcal{N}}\rho_{n}},$ (5)
where $l_{n}^{S}$ is the GPU consumption of the current task computed by ES. $c_{s}=o_{s}f_{s}$ denotes the GPU computation capability of ES. $o_{s}$ represents the FLOPS executed per cycle by the GPU of ES and $f_{s}$ denotes the GPU clock frequency of ES. The computational resources of the ES are evenly distributed among all users currently being served. Combining the aforementioned factors, the processing latency of UEn’s current task $q_{n}^{\mathcal{M}}$ can be expressed as
$t^{\mathcal{M}}_{n}=$ $\rho_{n}(t_{n}+t_{n}^{U}+t_{n}^{S})+(1-\rho_{n})t_{n}^{U}$ (6)
$=$ $\rho_{n}(t_{n}+t_{n}^{S})+t_{n}^{U}.$
@@ II-E Multi-Task Offloading Design and Problem Formulation
We consider three types of tasks: image classification, text classification, and VQA, all of which can be evaluated using accuracy to measure task performance:
@@
where $\varepsilon_{n}^{\mathcal{M}}$ is the accuracy of task $\mathcal{M}$, and $f^{\mathcal{M}}(\cdot)$ is the pretrained semantic transmission model for modality $\mathcal{M}$, and $\nu_{n}$ is SNR. Once the model training is completed, we can directly calculate the number of FLOPS required by the UE’s semantic encoder to determine the basic GPU consumption $l_{n}^{SE}$ for semantic extraction. Similarly, we compute the number of FLOPS required by the ES’s semantic decoder to determine the GPU consumption $l_{n}^{S}$ for the ES. To evaluate the performance of offloading tasks for different UEs, it is essential to define a unified QoE standard.
@@
The performance of task execution is associated with computation time, energy consumption, and task accuracy. However, defining a fair and reasonable QoE is challenging. When tasks of different modalities are unevenly distributed among different UEs, directly calculating the weighted sum of energy consumption, computation time, and task accuracy is unfair. Typically, text tasks have lower computation time and energy consumption than image tasks, while image tasks have lower computation time and energy consumption than VQA tasks. Additionally, the accuracies of different tasks are usually incomparable. Inspired by [41, 42, 43], we propose a QoE calculation method that can fairly compare task execution performance across different UEs:
$\mathrm{QoE}_{n}$ $=\sum_{q_{n}^{\mathcal{M}}=1}^{Q_{n}}(\omega_{t}G^{t}_{\mathcal{M}}+\omega_{e}G^{e}_{\mathcal{M}}+\omega_{a}G^{a}_{\mathcal{M}})$ (11)
@@
$\quad\left.+\frac{\omega_{a}}{1+e^{-\eta(\varepsilon_{n}^{\mathcal{M}}-\varepsilon^{l}_{n})}}\right),$
where $\omega_{t}$, $\omega_{e}$ and $\omega_{a}$ are the weights of computation time, energy consumption, and task accuracy, respectively. The sum of $\omega_{t}$, $\omega_{e}$ and $\omega_{a}$ are equal to 1. $t^{l}_{n}=l_{n}^{U}/c_{n}$ denotes the task execute delay by UEn. As shown in Fig. 3, we use the logistic/sigmoid function and its symmetrical counterpart for QoE calculations. Corresponding latency score $G^{t}_{\mathcal{M}}$ and energy consumption score $G^{e}_{\mathcal{M}}$ are calculated using the symmetric logistic function, while accuracy score $G^{a}_{\mathcal{M}}$ is measured with the logistic function. The offset $x_{0}$ represents local execution latency, energy consumption, or accuracy. Specifically, $E_{n}^{l}=\kappa^{U}t_{n}^{U}(f_{n})^{3}$ denotes the task execution energy consumption by UEn. $\varepsilon_{n}^{l}$ denotes the task executes accuracy by UEn. That is, we normalize the QoE using the latency, energy consumption, and accuracy of tasks executed locally by the UE.
@@
$n=1,2,\ldots,N,$ (12l)
where $\{\boldsymbol{\rho},\boldsymbol{p},\boldsymbol{f},\boldsymbol{x},\boldsymbol{\mu}\}=\{\rho_{n},p_{n},f_{n},x_{k,n},\mu_{n}\},\forall{n}\in\mathcal{N},\forall{k}\in\mathcal{K}$. $\boldsymbol{x}\in\mathbb{R}^{K\times N}$ denote the collection of all sub-channel allocation indicators $x_{k,n}$. $x_{k,n}=1$ indicates UEn is assigned to sub-channel $k$, $x_{k,n}=0$ otherwise. In (P0), (12b) and (12c) represent the offloading choice and channel selection constraints, respectively. Constraint (12d) and (12e) indicate that each UE is assigned to one sub-channel and each sub-channel is occupied by one UE. $p_{max}$ in (12g) is the maximum transmit power constraint, $f_{max}$ in (12h) is the maximum GPU frequency of UEn, $t_{max}$ in (12i) is the maximum execute latency constraint, $E_{n}^{max}$ in (12j) is the battery capacity of UEn, and $\varepsilon_{min}$ in (12k) is the minimum task accuracy requirement, which is dependent on the modality $\mathcal{M}$ of the current task.

### v2

@@ Abstract
Mobile edge computing (MEC) provides low-latency offloading solutions for computationally intensive tasks, effectively improving the computing efficiency and battery life of mobile devices. However, for data-intensive tasks or scenarios with limited uplink bandwidth, network congestion might occur due to massive simultaneous offloading nodes, increasing transmission latency, and affecting task performance. In this paper, we propose a semantic-aware multi-task offloading framework to address the challenges posed by limited uplink bandwidth. By introducing a semantic extraction factor, we balance the tradeoff among transmission latency, computation energy consumption, and task performance. To measure the offloading performance of different tasks, we design a unified and fair quality of experience (QoE) metric that includes execution latency, energy consumption, and task performance. We formulate the QoE optimization problem as a Markov decision process (MDP) and exploit the semantic-aware multi-agent proximal policy optimization (MAPPO) reinforcement learning algorithm to jointly optimize the semantic extraction factor and communication and computing resources allocation to maximize overall QoE. Experimental results show that the proposed method achieves a 12.68% improvement in user QoE compared with the semantic-unaware approach. Moreover, the proposed approach can be easily extended to models with different user preferences.
@@ I Introduction
Nevertheless, when tasks involve substantial data transfers, such as high-definition (HD) video processing [2], virtual reality (VR) applications [3], or real-time data analytics [4], the increased volume of data transmission imposes higher demands on channel quality to meet latency requirements. In such cases, semantic communication (SemCom) can provide an effective solution. SemCom, different from the traditional bit workflow based on Shannon’s information theory [5], has demonstrated great potential for enhancing the efficiency and reliability of data transmission across various task scenarios, including text [6, 7, 8], image [9, 10, 11], video [12, 13, 14], XR [15, 16, 17], and multi-task [18, 19, 20]. By transmitting a compact semantic representation of task data, substantial bandwidth can be saved, ensuring high quality of service (QoS).
@@ I-A Related Works
Given these considerations, it is essential to design a semantic-aware MEC system by combining MEC with SemCom [30]. By performing semantic extraction on the data to be offloaded, UEs can filter out data irrelevant to the task, thereby reducing network traffic. In [31], the authors defined a semantic entropy to quantify the semantic information of different tasks. Deep Q-Network (DQN) and many-to-one matching algorithms are used to solve the joint optimization problem of semantic compression, channel allocation, and power distribution in multi-cell multi-task scenarios. A semantic-aware resource allocation method was proposed in [32], introducing a semantic-aware factor to characterize the relationship between semantic compression rate and computational overhead, aiming to balance the degree of semantic compression. The authors of [33] proposed a semantic-aware cloud-edge-end collaborative networking method to utilize distributed computing resources and also proposed a multi-task offloading system in [34], where a unified QoE metric was designed.
@@ I-B Motivations and Contributions
In traditional MEC scenarios, computational and communication resources are often jointly optimized to minimize system energy consumption [35] or execution latency [36], or a combination of both [37]. However, these approaches lack consideration for task execution performance and comprehensive QoE. 6G strives to provide stable, reliable, and high-quality personalized services to each user [38]. Preferences often vary among different users and types of services. For instance, in MEC scenarios, users dealing with latency-sensitive tasks may prefer lower latency, those pursuing task performance may favor better task execution accuracy, while users with limited battery power might opt for task offloading strategies that consume less power. Therefore, a good QoE metric should encompass the preferences of different users. Additionally, since each user may need to handle different tasks, the complexity and evaluation metrics of different tasks vary. For example, bimodal tasks like visual question answering (VQA) typically require more processing time and computational resources compared to text translation, and their evaluation metrics are entirely different.
Semantic-aware MEC systems have been studied in [31, 32, 33, 34]. Nonetheless, in [31, 32, 33], the authors did not consider multi-tasks, limiting their ability to compare users engaged in different types of tasks. The authors of [34] did not consider the impact of semantic extraction on task execution latency and performance. To address the aforementioned challenges, we investigate resource allocation schemes in multi-task offloading that coordinate computational and communication resources through semantic compression-aware methods. To fairly compare the offloading performance of different task types, a unified QoE metric is proposed. Computational and communication resources are jointly optimized to maximize the overall QoE for UEs. The main contributions of this paper are summarized as follows:
  • We propose a novel semantic-aware task offloading system for multi-tasks. The SemCom systems for image classification, text classification, and VQA are first constructed to compress source data into compact semantic representations. The semantic extraction factor is used to characterize the degree of data compression, which affects the task’s execution latency, energy consumption, and performance.
@@
  • The joint optimization for semantic extraction factor, computational resources, and communication resources is performed using the semantic-aware MAPPO algorithm, maximizing the overall QoE. Users solve for the optimal channel allocation, power allocation, semantic extraction level, and computation offloading scheme according to their observed task sizes and channel conditions.
  • Comprehensive experiments demonstrate the effectiveness of the proposed approach. Compared to the semantic-unaware method, the proposed algorithm demonstrates a 12.68% improvement in user QoE. Additionally, the proposed approach can meet user preferences for various metrics such as execution latency, energy consumption, and execution performance.
@@ II System Model and Problem Formulation
We propose a novel semantic-aware multi-task offloading system, as shown in Fig. 1, where UEs employ an orthogonal frequency-division multiple access (OFDMA) scheme. At the access point (AP), $K$ resource blocks are allocated to serve $N$ users at different sub-channels. The set of UEs and sub-channels are defined as $\mathcal{N}=\{1,2,\ldots,N\}$ and $\mathcal{K}=\{1,2,\ldots,K\}$. The length of the UEn’s task queue is denoted as $Q_{n}$. The task currently being executed by the UEn is represented by $q^{\mathcal{M}}_{n}$, where $\mathcal{M}\in\{task_{1},task_{2},\ldots,task_{m}\}$. Hence, each UE may contain up to $m$ different types of tasks. We assume that tasks are atomic [34], i.e., all tasks are either executed locally or offloaded to the ES. The illustrations of the main notations are summarized in Table I.
@@ II-A Semantic-Aware Model
Traditional MEC offloading methods usually require transmitting all data to the ES. SemCom enables the extraction of task-relevant semantic information, reducing bandwidth consumption while enhancing robustness to channel variations. Based on this, we propose a semantic-aware multi-task offloading framework. Fig. 2 illustrates the process of semantic-aware offloading. While semantic extraction indeed reduces the volume of data transmitted, the associated increase in semantic extraction latency and computational burden must also be comprehensively considered. Hence, we define a semantic extraction factor $\mu_{n}\in\left[\mu_{n}^{min},1\right]$. The semantic extraction module determines the level of data compression. When $\mu_{n}=1$, the UE transmits all data instead of semantic information. Transmitting all data can achieve better task performance, but it imposes a heavy transmission burden. The smaller the $\mu_{n}$, the greater the level of semantic compression, which significantly reduces the transmission load. However, semantic extraction introduces additional computational overhead and may decrease task performance on the ES. The minimum value $\mu_{n}^{min}$ is typically determined by task accuracy requirements and user battery capacity. Therefore, selecting $\mu_{n}$ requires a comprehensive evaluation of the current task computational load, channel environment, user battery capacity, and other relevant information.
Figure 1: Semantic-aware multi-task offloading system model. Each task is fed into a semantic codec specifically designed for its type.
@@ II-B Transmission Model
The channel gain from UEn to ES via sub-channel $k$ is denoted by $|g_{k,n}|^{2}=|h_{k,n}|^{2}/d_{n}^{-\alpha}$, where $h_{n}\sim\mathcal{C}\mathcal{N}(0,1)$ is Rayleigh fading coefficient, and $d_{n}$ is the distance from UEn to ES, and $\alpha$ is the path loss exponent, and $\mathcal{C}\mathcal{N}(0,1)$ is the complexed Gaussian distribution with zero mean and variance one. The maximum achievable offloading data rate from UEn to ES can be represented as
$R_{n}=\sum_{k=1}^{K}B\text{log}_{2}\left(1+\frac{x_{k,n}\left|g_{k,n}\right|^{2}p_{n}}{\sigma_{z}^{2}}\right),$ (1)
where $B$ is the sub-channel bandwidth, and $p_{n}$ denotes the transmit power from UEn to ES, and $\sigma_{z}^{2}$ is the additive white Gaussian noise (AWGN) noise power. $x_{k,n}$ denotes the sub-channel allocation indicators. $x_{k,n}=1$ indicates UEn is assigned to sub-channel $k$, $x_{k,n}=0$ otherwise. Therefore, the transmission latency can be expressed as
$t_{n}=\frac{\rho_{n}s_{n}\mu_{n}^{p}}{R_{n}},$ (2)
where $s_{n}$ is the normalized transmission data from UEn to ES, and $p>0$ is a constant parameter. $\rho_{n}$ is an offload indicator, where $\rho_{n}=1$ represents offloading tasks to ES, and $\rho_{n}=0$ denotes local computation. Compared to the amount of data required for task offloading, the computation results typically involve a much smaller data size [39, 40], therefore, we only consider the upload rate and latency.
@@
$\mathcal{M}$ The set of task types
@@
$Q_{n}$ The task length of UEn
$q_{n}^{\mathcal{M}}$ The task currently being executed
@@ II-C Computation Model
Each task can be executed in one of three modes: locally on the device, directly offloaded to the ES ($\mu_{n}=1$), or semantically pre-processed locally and then offloaded ($\mu_{n}^{min}\leq\mu_{n}<1$). The latency incurred by semantic extraction at the user should be considered in the overall task execution time. Therefore, the computation latency of UEn can be denoted as
$t_{n}^{U}=\begin{cases}\frac{(1-\rho_{n})l_{n}^{U}}{c_{n}},\ \text{if}\ \mu_{n}=1,\\ \frac{(1-\rho_{n})l_{n}^{U}}{c_{n}}+\frac{\rho_{n}l_{n}^{SE}}{\mu_{n}^{q}c_{n}},\ \mu_{n}^{min}\leq\mu_{n}<1,\end{cases}$ (3)
where $q>0$ is a constant parameter, $l_{n}^{U}$ is the Graphic Processing Unit (GPU) consumption of the current task computed by UEn, and $l_{n}^{SE}$ is the basic GPU consumption of semantic extraction. $l_{n}^{U}$ and $l_{n}^{SE}$ are related to the corresponding task $\mathcal{M}$. $k>0$ is a constant parameter. $c_{n}=\alpha_{ma}N_{U}f_{n}$ is the GPU computation capability of UEn. $\alpha_{ma}=2$ corresponds to the floating-point operations (FLOPs) per cycle per core (FLOPs/cycle/core), where each instruction executes a fused multiply-add (FMA) operation. $N_{u}$ represents the number of CUDA cores available on UEn. $f_{n}$ denotes the GPU clock frequency of UEn. When tasks are offloaded to ES, the computation time of tasks on the ES can be expressed as
$t^{S}=\frac{\sum_{n=1}^{N}l_{n}^{S}}{c_{s}},$ (4)
where $c_{s}=N_{S}f_{s}\alpha_{ma}$ is the GPU computation capability of ES. $l_{n}^{S}$ is the GPU consumption of the task offloaded by the UEn to the ES. $N_{s}$ denotes the number of CUDA cores available on ES. $f_{s}$ denotes the GPU clock frequency of ES. ES allocates its computational resources among users fairly, proportional to the amount of computation offloaded by each user. Combining the aforementioned factors, the processing latency of UEn’s current task $q_{n}^{\mathcal{M}}$ can be expressed as
$t^{\mathcal{M}}_{n}=$ $\rho_{n}(t_{n}+t_{n}^{U}+t^{S})+(1-\rho_{n})t_{n}^{U}$ (5)
$=$ $\rho_{n}(t_{n}+t^{S})+t_{n}^{U}.$
@@ II-E Multi-Task Offloading Design and Problem Formulation
The execution performance of all tasks in the task queue is represented as
@@
where $\varepsilon_{n}^{\mathcal{M}}$ is the accuracy of task $\mathcal{M}$, and $f^{\mathcal{M}}(\cdot)$ is the pretrained semantic transmission model for task $\mathcal{M}$, and $\nu_{n}$ is SNR. Once the model training is completed, we can directly calculate the number of FLOPs required by the UE’s semantic encoder to determine the basic GPU consumption $l_{n}^{SE}$ for semantic extraction. Similarly, we compute the number of FLOPs required by the ES’s semantic decoder to determine the GPU consumption $l_{n}^{S}$ for the ES. To evaluate the performance of offloading tasks for different UEs, it is essential to define a unified QoE standard.
@@
The performance of task execution is associated with computation time, energy consumption, and task accuracy. However, defining a fair and reasonable QoE is challenging. When tasks of different types are unevenly distributed among different UEs, directly calculating the weighted sum of energy consumption, computation time, and task accuracy is unfair. Typically, text tasks have lower computation time and energy consumption than image tasks, while image tasks have lower computation time and energy consumption than VQA tasks. Additionally, the accuracies of different tasks are usually incomparable. Inspired by [41, 42, 43], we propose a QoE calculation method that can fairly compare task execution performance across different UEs:
$\mathrm{QoE}_{n}$ $=\frac{1}{Q_{n}}\sum_{q_{n}^{\mathcal{M}}=1}^{Q_{n}}(\omega_{t}G^{t}_{\mathcal{M}}+\omega_{e}G^{e}_{\mathcal{M}}+\omega_{a}G^{a}_{\mathcal{M}})$ (10)
@@
$\left.+\frac{\omega_{a}}{1+e^{-\eta(\varepsilon_{n}^{\mathcal{M}}-\varepsilon^{l}_{n})}}\right),$
where $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are the weights of computation time, energy consumption, and task accuracy, respectively. The sum of $\omega_{t}$, $\omega_{e}$ and $\omega_{a}$ are equal to 1. $Q_{n}$ is the task length of UEn. $t^{l}_{n}=l_{n}^{U}/c_{n}$ denotes the task execution delay by UEn. As shown in Fig. 3, we use the logistic/sigmoid function and its symmetrical counterpart for QoE calculations. Corresponding latency score $G^{t}_{\mathcal{M}}$ and energy consumption score $G^{e}_{\mathcal{M}}$ are calculated using the symmetric logistic function, while accuracy score $G^{a}_{\mathcal{M}}$ is measured with the logistic function. The offset $x_{0}$ represents local execution latency, energy consumption, or accuracy. Specifically, $E_{n}^{l}=\kappa^{U}t_{n}^{l}(f_{n})^{3}$ denotes the task execution energy consumption by UEn. $\varepsilon_{n}^{l}$ denotes the task execution accuracy by UEn. That is, we normalize the QoE using the latency, energy consumption, and accuracy of tasks executed locally by the UE.
@@
where $\{\boldsymbol{\rho},\boldsymbol{p},\boldsymbol{f},\boldsymbol{x},\boldsymbol{\mu}\}=\{\rho_{n},p_{n},f_{n},x_{k,n},\mu_{n}\},\forall{n}\in\mathcal{N},\forall{k}\in\mathcal{K}$. $\boldsymbol{x}\in\mathbb{R}^{K\times N}$ denote the collection of all sub-channel allocation indicators $x_{k,n}$. $x_{k,n}=1$ indicates UEn is assigned to sub-channel $k$, $x_{k,n}=0$ otherwise. In (P0), (11b) and (11c) represent the offloading choice and channel selection constraints, respectively. Constraints (11d) and (11e) indicate that each UE is assigned to one sub-channel and each sub-channel is occupied by one UE. $p_{max}$ in (11g) is the maximum transmit power constraint, $f_{max}$ in (11h) is the maximum GPU frequency of UEn, $t_{max}$ in (11i) is the maximum execute latency constraint, $E_{n}^{max}$ in (11j) is the battery capacity of UEn, and $\varepsilon_{min}$ in (11k) is the minimum task accuracy requirement, which is dependent on the task type $\mathcal{M}$ of the current task.